OpenStudy (anonymous):

Choose the equation of the line passing through the point (–2, 6) and parallel to y = 1/2x – 8. y = 1/2x + 7 y = 1/2x + 5 y = 2x + 7 y = 2x + 5

OpenStudy (anonymous):

Use same slope, so y = (1/2)x + b Now, substitute the point: 6 = (1/2)(-2) + b, so b = 7

OpenStudy (anonymous):

thx for the recognition! Good luck in all of your studies!